OpenStudy (anonymous):

Solve: 2(x + 1) = 2x + 5 Any Real Number No Solution 0 3

OpenStudy (anonymous):

2x+2=2x+5 0x=3 0=/=3 No solution

OpenStudy (anonymous):

this is like the last one, first multiply the first one out

OpenStudy (anonymous):

cuz that'll teach her >.>

OpenStudy (anonymous):

multiply 2(x + 1) out 2 times x= 2x 2 times 1= 2 now you have 2x+1 = 2x+5

OpenStudy (anonymous):

since there is a 2x on each side you can cancle those out and have 1=5 since we learned in first grade that 1 does NOT equal 5 then there is no solution